OpenStudy (anonymous):

Use synthetic division and the factor theorem to match the polynomial with its correct factor. x^3 + x^2 + 13x +4 x^3 - x +6 x^3 + x^2 - 3x +9

OpenStudy (anonymous):

answers x + 4 x + 2 x+3 no factor listed for this polynomial

OpenStudy (anonymous):

this is a lot of donkey work you want to do it the snap way?

OpenStudy (anonymous):

\(x^3 + x^2 + 13x +4 \) doesn't have any factors with integers

OpenStudy (anonymous):

sorry its just that this is my last work in its like 20 questions

OpenStudy (anonymous):

in i want to get it done so i can be over with it

OpenStudy (anonymous):

i understand i am just saying that this requires several steps of synthetic division, which is easy but a pain so it is easier to factor each polynomial

OpenStudy (anonymous):

\[x^3-x+6=(x+2) (x^2-2 x+3)\]so you see the factor right away

OpenStudy (anonymous):

ok im lost

OpenStudy (anonymous):

you are asked to match the polynomial with its factor

OpenStudy (anonymous):

hold on i screwed up sorry

OpenStudy (anonymous):

i meant since \(x^3-x+6=(x+2) (x^2-2 x+3)\) you know that \(x+2\) is a factor of \(x^3-x+6\)

OpenStudy (anonymous):

so you can match up \(x+2\) with \(x^3-x+6\)

OpenStudy (anonymous):

it is a weird way to ask the question

OpenStudy (anonymous):

similarly \(x^3 + x^2 - 3x +9 =(x+3) (x^2-2 x+3)\) so you can match up \(x+3\) with \(x^3 + x^2 - 3x +9 \)

OpenStudy (anonymous):

that is all you can do for this one

OpenStudy (anonymous):

IN THEN x^3 + x^2 + 13x +4 FOR no factor listed for this polynomial

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

let me make sure hold on one second

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

here you can see for yourself that it does not factor http://www.wolframalpha.com/input/?i=x^3+%2B+x^2+%2B+13x+%2B4+

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw 19 to go

OpenStudy (anonymous):

Match the polynomial with its degree or the function with its value f(x) = x^2 - 3x + 3 g(x) = 2x^2 -1 g(3) f(z^2) f(4) x^4 + 3x^2 -5x 5x^10 + 13x^4 -17x^3 g(w^3) f(z) 5x^2 - 3x^5 -5x^3

OpenStudy (anonymous):

Choose... 7 2w6 -1 z4 -3z2 +3 4 17 z2 -3z +3 10 5

OpenStudy (anonymous):

can u help!!